OpenStudy (anonymous):

A stop sign is a regular octagon. This stop sign has a height of 30 inches and each side is 12.43 inches. The area of a 30-inch sign is ____ in2 to the nearest square inch. (Enter only the number.)

OpenStudy (anonymous):



OpenStudy (whpalmer4):

Do you know a formula for the area of a regular octagon?

OpenStudy (anonymous):

Area of a octagon =

OpenStudy (anonymous):

Oh it didn't... hold on ahah

OpenStudy (anonymous):

Let S = the length of the side of the octagon. Let A = the area of the octagon Then A = 2(1+ sqrt2)xS^2 = 4.828 S^2

OpenStudy (whpalmer4):

You can look up such a formula, or we can figure it out. Because this is a regular octagon, we can divide it up into a bunch of isoceles triangles with their apex at the center, and the base being the outside edges.

OpenStudy (whpalmer4):

That'd be great, if only we knew the side length of the octagon. Do you see a way to figure it out?

OpenStudy (anonymous):

Lol it says in the question 12.43

OpenStudy (whpalmer4):

Great! That's the easiest way to find it :-)

OpenStudy (whpalmer4):

What's your answer for the area?

OpenStudy (anonymous):

I have to multiply 12.43 times 12.43 right?

OpenStudy (whpalmer4):

Yes. You'll be evaluating \[A=2(1+\sqrt{2})(S)^2 \approx 4.828*12.43*12.43 =\]

OpenStudy (anonymous):

12.43x12.43=154.5049

OpenStudy (whpalmer4):

Agreed

OpenStudy (whpalmer4):

But that's not the final answer...

OpenStudy (anonymous):

I know...I suck at math I need my hand held in it Lol

OpenStudy (whpalmer4):

\[A = 4.828*12.43*12.43 = 4.828*154.5049 =\]

OpenStudy (anonymous):

so 4.828x154.5049 and the sum of that is the answer?

OpenStudy (whpalmer4):

Multiply 4.828 * 154.5049. The number you get (the product), rounded to the nearest square inch, is the answer.

OpenStudy (whpalmer4):

The 4.828 replaces the \(2(1+\sqrt{2})\) in the formula.

OpenStudy (anonymous):

Okay so I got 745.9496572

OpenStudy (whpalmer4):

Yes, that's the product of those numbers, but not your answer... "rounded to the nearest square inch"

OpenStudy (anonymous):

Haha I love your patience. As you can see I suck at math I'm like a 5 year old... How do you "round to the nearest square inch" .__. I feel dumb haha :P

OpenStudy (whpalmer4):

Okay, to round a number, you look at the next digit after the spot you are rounding to. If it is 5 or larger, you round your number up and zero out everything to the right. If it is 0-4, you just zero out everything to the right. Some examples: 53 rounded to the nearest 10: Look at the 3 in the 1s place, that's less than 5, so our answer is 50. 57 rounded to the nearest 10: Look at the 7 in the 1s place, that's 5 or greater, so our answer is 60. 2.41 rounded to the nearest 0.1: Look at the 1 in the .01 place, that's less than 5, so our answer is 2.4 2.45 rounded to the nearest 0.1: Look at the 5 in the .01 place, that's 5 or greater, so our answer is 2.5 Does that make sense? Prove it by rounding off the answer correctly :-)

OpenStudy (anonymous):

745.90?

OpenStudy (whpalmer4):

No. Isn't the nearest square inch going to be a nice round number like 756 or 835, and not have a decimal after it? How do you read that number? Seven hundred forty-five and 90 one-hundredths. Sounds like there's a fraction in there, doesn't it?